# New Applications of Ergodic Theory to Sets of Differences

### v1

@@ 3.2 Proof of a General Variant of Theorem 1
Remark 3.11. Dealing with sets of the form $\Delta_{1,g}(S)$ brings to life new phenomena which distinguish Theorem 3.2 from Theorem 1. For example, let $g$ be a function coming from a Hardy field with the property that for some $\ell\in\mathbb{N}\cup\{0\}$, $\lim_{x\to\infty}g^{(\ell)}(x)=\pm\infty$ and $\lim_{x\to\infty}g^{(\ell+1)}(x)=0$. If $([g(n)])_{n=1}^{\infty}$ is pointwise ergodic then it follows from [9, Theorem A] that $D_{g}$ is thick (and not necessarily syndetic). For instance, if $g$ is one of the functions $n\log(n)$, $\frac{n^{2}}{\log(n)}$, or $n^{c}$ for $c>0$ with $c\not\in\mathbb{N}$, which are listed in (3.9), the set $D_{g}$ is thick. On the other hand one can show, for a large class of Hardy functions which satisfy $\lim_{n\to\infty}g^{(\ell)}(n)\to K$ for some finite constant $K>0$ the set $D_{g}$ is syndetic. This is the case for $g(n)=\sqrt{2}n^{2}+\sqrt{3}n$.
@@ 4 Generalizing Theorem 1 to the Amenable Setup
4.1 Amplifying Theorem 1
The ergodic techniques that were used in the previous sections admit further amplifications, which will allow us to establish a rather general form of Theorem 1 for countably infinite amenable groups and cancellative semigroups. Actually as we will see in Theorem 4.1 below, the amenable point of view allows for interesting extensions of Theorem 1 even for subsets of $\mathbb{Z}$. Indeed, a careful examination of the proof of Theorem 1 in Section 2 reveals that the assumption that the sets $S_{1},\ldots,S_{k}$ have positive upper density can be significantly weakened.
Before formulating Theorem 4.1, we need to introduce some additional notation. Recall that a Følner sequence in $\mathbb{Z}$ is a sequence $(F_{N})_{N=1}^{\infty}$ of finite nonempty subsets of $\mathbb{Z}$ satisfying
@@ 4.1 Amplifying Theorem 1
Let $\mathbf{F}=(F_{N})$ be a Følner sequence in $\mathbb{Z}$. Given a set $S\subseteq\mathbb{Z}$, define
$\overline{d}_{\mathbf{F}}(S)=\limsup_{N\rightarrow\infty}\frac{|S\cap F_{N}|}{|F_{N}|},$ (4.2)
if the limit in formula (4.2) exists then its value is called the density of $S$ along $\mathbf{F}$ and is denoted $d_{\mathbf{F}}(S)$. Additionally, let
@@
It is worth mentioning that it is often natural to consider subsets and Følner sequences in $\mathbb{N}$; the above discussion equally applies to this case.
Note that $\Delta_{2}(\mathbf{F},S)$ coincides with $\Delta_{2}(S)$ and $\overline{d}_{\mathbf{F}}(S)$ coincides with $\overline{d}(S)$ as defined in Section 1 when $\mathbf{F}=(F_{N})$ is the standard Følner sequence $F_{N}=\{1,\ldots,N\}$.
@@
Theorem 4.1. Let $\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ be Følner sequences in $\mathbb{Z}$. If $S_{1},\ldots,S_{k}\subseteq\mathbb{Z}$ satisfy $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$, then there exists $S\subseteq\mathbb{Z}$ such that $d(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S)\subseteq D=\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in\mathbb{Z}$, where
@@
One can also obtain an amplification of Theorem 3.2 similar to Theorem 4.1. Before stating the amplified version of Theorem 3.2, we recall/introduce some pertinent notation. For any function $g:\mathbb{N}\rightarrow[1,\infty)$ and any Følner sequence $\mathbf{F}$ in $\mathbb{N}$, let
@@
Theorem 4.2. Let $g:\mathbb{N}\rightarrow[1,\infty)$ and assume $([g(n)])_{n=1}^{\infty}$ is pointwise ergodic. Let $\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ be Følner sequences in $\mathbb{N}$. If $S_{1},\ldots,S_{k}\subseteq\mathbb{N}$ satisfy $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$, then there exists $S\subseteq\mathbb{N}$ such that $d(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1,g}(S)\subseteq D_{g}:=\bigcap_{i=1}^{k}\Delta_{2,g}(\mathbf{F}_{i},S_{i})$. Furthermore, we have that $\overline{d}(D_{g})\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$.
The proofs of Theorems 4.1 and 4.1 are practically the same as those of Theorems 1 and 3.2, respectively, with one modification which involves the use of a more general version of Theorem 2 in which $\overline{d}$ is replaced by $\overline{d}_{\mathbf{F}}$ (this result is a special case of Theorem 4.2 which is formulated in the next subsection).
Note that in Theorem 4.1, besides the Følner sequences $\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$, one more Følner sequence is implicitly present. Namely, $d(S)$ actually means $d_{\mathbf{F}}(S)$, where $\mathbf{F}$ is the standard Følner sequence $\mathbf{F}=(F_{N})$, $F_{N}=\{1,\ldots,N\}$. The role of the Følner sequence $\mathbf{F}$ is different from the roles of the sequences $\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$. While $\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ are needed just to define appropriate notions of largeness for $S_{1},\ldots,S_{k}$, the role of the sequence $\mathbf{F}$ is to guarantee that the ergodic averages converge pointwise (which is used in the proof of Theorem 2 in [3]). These ideas will be clarified in the next subsection where we extend Theorem 4.1 to the general amenable setup.
@@ 4.2 Extending Theorem 1 to amenable groups
Let $G$ be a countably infinite semigroup. A left Følner sequence in $G$ is a sequence $(F_{N})$ of finite nonempty subsets of $G$ satisfying
@@
for all $g\in G$. It is well-known that the existence of a left Følner sequence in a countably infinite semigroup $G$ implies that $G$ is left amenable, furthermore if $G$ is cancellative then having a left Følner sequence is equivalent to $G$ being left amenable. From now on, we will tacitly assume that the semigroups we deal with are countable and cancellative.
Given a left Følner sequence $\mathbf{F}=(F_{N})$ in $G$ and a set $S\subseteq G$, define
@@
If $\overline{d}_{\mathbf{F}}(S)=\underline{d}_{\mathbf{F}}(S)$ then their common value is denoted by $d_{\mathbf{F}}(S)$.
Definition 4.3. For a group $G$, a sequence $(F_{n})_{n=1}^{\infty}$ of subsets of $G$ is called left tempered if there is some constant $C>0$ such that
@@
[…]
From now on we will, as a rule, omit the adjective left when dealing with left amenable semigroups and left Følner sequences.
In order prove the generalization of Theorem 1 to countably infinite amenable groups we will need two results which can be viewed as amenable analogs of Theorems 2 and 2.
Theorem 4.4 (Theorem 2 for countably infinite amenable groups). Let $G$ be a countably infinite amenable group with a tempered Følner sequence $\mathbf{F}$. Let $(X,\mathcal{B},\mu,(T_{g})_{g\in G})$ be a measure-preserving system and let $B\in\mathcal{B}$ with $\mu(B)>0$. Then there is a set $P\subseteq G$ such that
$d_{\textbf{F}}(P)\geq\mu(B)\quad\text{and}\quad\mu\left(B\cap T_{g_{1}}^{-1}B\cap\cdots\cap T_{g_{r}}^{-1}B\right)>0\quad\text{for all}\quad g_{1},...,g_{r}\in P$ (4.8)
Proof. After possibly subtracting a set of measure zero from $X$, we may assume, without loss of generality, that if $g_{1},\ldots,g_{r}\in G$ and $T_{g_{1}}^{-1}B\cap\cdots\cap T_{g_{r}}^{-1}B\neq\emptyset$ then
$\mu\left(T_{g_{1}}^{-1}B\cap\cdots\cap T_{g_{r}}^{-1}B\right)>0$. Denote $\mathbf{F}$ by $(F_{N})_{N=1}^{\infty}$. For $N=1,2,3,\ldots$ define $f_{N}:X\rightarrow\mathbb{R}$ by,
@@
For all $N$ we have $\int f_{N}=\mu(B)$. Let $f=\limsup_{N}f_{N}$. Note that $0\leq f_{N}\leq 1$ for all $N$, so $f$ is well-defined and takes values in $[0,1]$. Then by Fatou’s Lemma we have
$\int f=\int\limsup_{N}f_{N}\geq\limsup_{N}\int f_{N}=\mu(B).$
Since $f$ is nonnegative and $\int f\geq\mu(B)$, the set $f^{-1}[\mu(B),1]$ must have positive measure. By a theorem from A. Shulman [16], $\lim_{N}f_{N}$ exists a.e. (for a precise account of Shulman’s result see Section 5.6 in [18]), so in particular there is some $x_{0}\in f^{-1}[\mu(B),1]$ such that $\lim_{N}f_{N}(x_{0})=f(x_{0})$. If we set $P^{\prime}\vcentcolon=\{g\in G:T_{g}x_{0}\in B\}$ then
$d_{\mathbf{F}}(P^{\prime})=\lim_{N}\frac{|F_{N}\cap P^{\prime}|}{|F_{N}|}=\lim_{N}\frac{1}{|F_{N}|}\sum_{g\in F_{N}}1_{B}({T_{g}}x_{0})=\lim_{N}f_{N}(x_{0})=f(x_{0})\geq\mu(B).$
Now choose $h\in G$ so that $T_{h}x_{0}\in B$ and put $P=P^{\prime}h^{-1}$. We claim that $P$ satisfies (4.8). Firstly, observe $d_{\mathbf{F}}(P)=d_{\mathbf{F}}(P^{\prime})\geq\mu(B)$. Secondly, take $g_{1}h^{-1},\ldots,g_{r}h^{-1}\in P$ where $g_{1},\ldots,g_{r}\in P^{\prime}$. See that for $i=1,\ldots,r$ we have $T_{g_{i}}x_{0}\in B$, so $T_{g_{i}h^{-1}}T_{h}x_{0}\in B$, therefore $T_{h}x_{0}\in T_{g_{i}h^{-1}}^{-1}B$. Thus the intersection $B\cap T_{g_{1}h^{-1}}^{-1}B\cap\cdots\cap T_{g_{r}h^{-1}}^{-1}B$ is nonempty since it contains $T_{h}x_{0}$, so $\mu\left(B\cap T_{g_{1}h^{-1}}^{-1}B\cap\cdots\cap T_{g_{r}h^{-1}}^{-1}B\right)>0$. ∎
For the proof of Theorem 4.2 we will need the following variant of Furstenberg correspondence in the amenable setup.
Theorem 4.5 ([5, Theorem 6.4.17]). Let $G$ be a countably infinite amenable group, let $\mathbf{F}$ be a Følner sequence in $G$, and let $S\subseteq G$ with $\overline{d}_{\mathbf{F}}(S)>0$. Then there is an invertible measure-preserving system $(X,\mathcal{B},\mu,(T_{g})_{g\in G})$ and a set $B\in\mathcal{B}$ with $\mu(B)=\overline{d}_{\mathbf{F}}(S)$ such that

### v2

@@ 3.1 Ergodic sequences
An interesting feature of Theorems 3.2 and 3.2 is that they provide sufficient conditions for sets of the form $D_{g}$ to be thick or syndetic. In particular, if the function $g$ comes from a Hardy field with the property that for some $\ell\in\mathbb{N}\cup\{0\}$, $\lim_{x\to\infty}g^{(\ell)}(x)=\pm\infty$ and $\lim_{x\to\infty}g^{(\ell+1)}(x)=0$, then the set $D_{g}$ is always thick, as mentioned in Theorem 3.2. Such is the case if $g$ is one of the functions $n\log(n)$, $\frac{n^{2}}{\log(n)}$, or $n^{c}$ for $c>0$ with $c\not\in\mathbb{N}$, which are listed in (3.9). On the other hand, if $g$ is a Hardy field function such that $([g(n)])_{n=1}^{\infty}$ is pointwise ergodic and $g$ satisfies $\lim_{x\to\infty}g^{(\ell)}(x)\to K$ for some finite constant $K\neq 0$, the set $D_{g}$ in Theorem 3.2 may be syndetic. For example, if $g(x)=a_{n}x^{n}+a_{n-1}x^{n-1}+\cdots+a_{0}\in\mathbb{R}[x]$ has at least two rationally independent coefficients other than the constant term, then $D_{g}$ is always syndetic (see Theorem 3.2). An example of a function $g:\mathbb{N}\to[1,\infty)$ (which satisfies $\lim_{x\to\infty}g^{\prime}(x)\to 2$) and a set $S\subseteq\mathbb{N}$ for which the set $D_{g}=\Delta_{2,g}(S)$ from Theorem 3.2 need not be thick nor syndetic is provided below by Example 3.3.
Example 3.13. If $g=2n+2\sqrt{n}$, then $([g(n)])_{n=1}^{\infty}$ is pointwise ergodic and the set $D_{g}=\Delta_{2,g}(4\mathbb{N})$ is neither thick nor syndetic.
Proof. It follows from [12, Theorem 3.4] that $([g(n)])_{n=1}^{\infty}$ is pointwise ergodic. We have
$D_{g}=\{n\in\mathbb{N}:\overline{d}(4\mathbb{N}\cap(4\mathbb{N}-[g(n)]))>0\}=\{n:[g(n)]\in 4\mathbb{N}\}$
Note that for any real number $t$ one has that $[t]\in 4\mathbb{N}$ if and only if $\{\frac{1}{4}t\}\in[0,\frac{1}{4})$. So,
$D_{g}=\left\{n:\left\{\frac{1}{4}g(n)\right\}\in\left[0,\frac{1}{4}\right)\right\}=\left\{n:\left\{\frac{1}{2}(n+\sqrt{n})\right\}\in\left[0,\frac{1}{4}\right)\right\}.$ (3.26)
To see that this set is not thick, observe that if $n\in D_{g}$ and $\epsilon=\frac{1}{2}(\sqrt{n+1}-\sqrt{n})<\frac{1}{4}$, then
$\left\{\frac{1}{2}((n+1)+\sqrt{n+1})\right\}=\left\{\frac{1}{2}+\frac{1}{2}(\sqrt{n+1}-\sqrt{n})+\frac{1}{2}(n+\sqrt{n})\right\}\in\left(\frac{1}{2},\frac{3}{4}+\epsilon\right).$ (3.27)
So we have shown that if $n\in D_{g}$, then $(n+1)\not\in D_{g}$, which implies $D_{g}$ is not thick.
Now we will show that $D_{g}$ is not syndetic. Fix $0<\epsilon<\frac{1}{8}$, then there exist arbitrarily large $n\in\mathbb{N}$ such that $\left\{\frac{1}{2}(n+\sqrt{n})\right\}\in\left(\frac{1}{4},\frac{1}{4}+\epsilon\right)$. Now observe for any $k\in\mathbb{N}$ and for $n\in\mathbb{N}$ large enough such that $\left\{\frac{1}{2}(n+\sqrt{n})\right\}\in\left(\frac{1}{4},\frac{1}{4}+\epsilon\right)$ we have for all $0\leq i\leq k$ that $\left\{\frac{1}{2}(n+i+\sqrt{n+i})\right\}\in\left(\frac{1}{4},\frac{1}{4}+2\epsilon\right)\cup\left(\frac{3}{4},\frac{3}{4}+2\epsilon\right)$. This means for any $k\in\mathbb{N}$ we can find a set $\{n,n+1,\dots,n+k\}\subset A^{c}$, hence $A$ is not syndetic. $\square$
The following example demonstrates that the set $D_{g}$ in Theorem 3.2, while guaranteed to be thick, need not be syndetic.
Example 3.14. Let $D_{3/2}=\Delta_{2,{3/2}}(2\mathbb{N})=\{n\in\mathbb{N}:\overline{d}(2\mathbb{N}\cap(2\mathbb{N}-[n^{3/2}]))>0\}$. Then $D_{3/2}$ is not syndetic.
Proof. Note that $D_{3/2}=\{n\in\mathbb{N}:[n^{3/2}]\in 2\mathbb{N}\}=\{n\in\mathbb{N}:\{\frac{1}{2}n^{3/2}\}\in[0,1/2)\}$. In order to prove $D_{3/2}$ is not syndetic, let $N\in\mathbb{N}$ and observe that it suffices to show that there is $M\in\mathbb{N}$ for which $M,M+1,\ldots,M+N\notin D_{3/2}$, which is equivalent to showing that $\{\frac{1}{2}(M+k)^{1/2}\}\in[1/2,1)$ for $k=0,1,\ldots,N$. To this end, we first observe the following:
  • (a)
    $\left(\frac{1}{2}(x+1)^{3/2}-\frac{1}{2}x^{3/2}\right)-\frac{3}{4}\sqrt{x}\rightarrow 0$ as $x\rightarrow\infty$,
  • (b)
    The sequence ($\left(\left\{\frac{1}{2}n^{3/2}\right\},\left\{\frac{3}{4}\sqrt{n}\right\}\right)$, $n\in\mathbb{N}$ is dense in $[0,1]^{2}$ (this sequence is actually uniformly distributed in $[0,1]^{2}$ – this follows from Féjer’s Theorem [21, Corollary 2.1, page 14] and the Weyl criterion [21, Theorem 6.2, page 48]),
  • (c)
    $\sqrt{x+1}-\sqrt{x}\rightarrow 0$ as $x\rightarrow\infty$ which, along with observation (b), implies that for any interval $(a,b)\subseteq[0,1]$, there exist arbitrarily large $m\in\mathbb{N}$ for which $\left\{\frac{3}{4}\sqrt{m+k}\right\}\in(a,b)$ for $k=0,1,...,N$.
Now, fix $\alpha<\beta$ such that $[\alpha,\beta]\subseteq(0,\frac{1}{16})$, and choose $\delta>0$ so that $(\frac{\alpha}{N}-\delta,\beta+\delta)\subseteq(0,\frac{1}{16})$. In light of the above observations, we can choose $M\in\mathbb{N}$ satisfying
$\begin{split}\text{(i)}\,\,\,\left\{\frac{1}{2}M^{3/2}\right\}\in\left(\frac{5}{8},\frac{7}{8}\right),\quad\text{(ii)}\,\,\,\left\{\frac{3}{4}\sqrt{M+j}\right\}\in\left(\frac{\alpha}{N},\frac{\beta}{N}\right)\text{ for }j=0,1,...,N,\text{ and }\\ \text{(iii)}\,\,\,\left(\frac{1}{2}(M+j)^{3/2}-\frac{1}{2}(M+j-1)^{3/2}\right)-\frac{3}{4}\sqrt{M+j-1}<\frac{\delta}{N}\text{ for }j=1,\ldots,N.\end{split}$ (3.28)
Let $1\leq k\leq N$ and recall that it suffices to show $\left\{\frac{1}{2}(M+k)^{3/2}\right\}\in[1/2,1)$. Observe that
$\left|\left(\frac{1}{2}(M+k)^{3/2}-\frac{1}{2}M^{3/2}\right)-\sum_{i=1}^{k}\frac{3}{4}\sqrt{M+i-1}\right|$
$=\left|\sum_{i=1}^{k}\left(\frac{1}{2}(M+i)^{3/2}-\frac{1}{2}(M+i-1)^{3/2}\right)-\sum_{i=1}^{k}\frac{3}{4}\sqrt{M+i-1}\right|$
$\leq\sum_{i=1}^{k}\left|\left(\frac{1}{2}(M+i)^{3/2}-\frac{1}{2}(M+i-1)^{3/2}\right)-\frac{3}{4}\sqrt{M+i-1}\right|<k\cdot\frac{\delta}{N}\leq N\cdot\frac{\delta}{N}=\delta.$ (3.29)
By (ii) in formula (3.28), it follows that $\left\{\sum_{i=1}^{k}\frac{3}{4}\sqrt{M+i-1}\right\}\in(\frac{\alpha}{N},\beta)$, and so by formula (3.29), we get that
$\left\{\frac{1}{2}(M+k)^{3/2}-\frac{1}{2}M^{3/2}\right\}=\left\{\left\{\frac{1}{2}(M+k)^{3/2}\right\}-\left\{\frac{1}{2}M^{3/2}\right\}\right\}\in\left(\frac{\alpha}{N}-\delta,\beta+\delta\right)\subseteq\left(0,\frac{1}{16}\right).$ (3.30)
Since fractional parts lie in $[0,1)$, it follows from formula (3.30) that either
[…]
By (i) from formula (3.28), $\{\frac{1}{2}(M+k)^{3/2}\}-\{\frac{1}{2}M^{3/2}\}>0-\frac{7}{8}>-\frac{15}{16}$, so case (1) is not possible. Hence, case (2) must hold and by (i) from formula (3.28), so it follows that $\{\frac{1}{2}(M+k)^{3/2}\}\in(\frac{10}{16},\frac{15}{16})\subseteq[\frac{1}{2},1)$ as desired. $\square$
@@ 4 Generalizing Theorem 1 to the Amenable Setup
The goal of this section is to develop some far-reaching extensions of Theorem 1. In Section 4.1 we obtain an amplification of Theorem 1 which applies to general families of “large sets” $S_{1},\dots,S_{k}\subset\mathbb{Z}$. In Section 4.2 we generalize Theorem 1 to the setup of countable amenable groups via Theorem 4.2. Finally, in Section 4.3 we generalize Theorem 4.2 to countable amenable cancellative semigroups.
4.1 Amplifications of Theorem 1 in $\mathbb{Z}$
The ergodic techniques that were used in the previous sections admit further amplifications, which will allow us to establish a rather general form of Theorem 1 for countably infinite amenable groups and cancellative semigroups. Actually, as we will see in Theorem 4.1 below, the point of view based on the notion of amenability allows for interesting extensions of Theorem 1 even for subsets of $\mathbb{Z}$. Indeed, a careful examination of the proof of Theorem 1 in Section 2 reveals that the assumption that the sets $S_{1},\ldots,S_{k}$ have positive upper density can be significantly weakened.
Before formulating Theorem 4.1, we need to introduce some additional notation. Recall that a Følner sequence in $\mathbb{Z}$ is a sequence $(F_{N})_{N=1}^{\infty}$ of finite non-empty subsets of $\mathbb{Z}$ satisfying
@@ 4.1 Amplifications of Theorem 1 in $\mathbb{Z}$
Let $\mathbf{F}=(F_{N})_{N=1}^{\infty}$ be a Følner sequence in $\mathbb{Z}$. Given a set $S\subseteq\mathbb{Z}$, define
$\overline{d}_{\mathbf{F}}(S)=\limsup_{N\rightarrow\infty}\frac{|S\cap F_{N}|}{|F_{N}|}.$ (4.2)
If the limit in formula (4.2) exists, then its value is called the density of $S$ along $\mathbf{F}$ and is denoted by $d_{\mathbf{F}}(S)$. Additionally, let
@@
It is worth mentioning that it is often natural to consider subsets and Følner sequences in $\mathbb{N}$ instead of $\mathbb{Z}$; the above definitions admit trivial modifications which are applicable to this case.
Note that $\Delta_{2}(\mathbf{F},S)$ coincides with $\Delta_{2}(S)$ and $\overline{d}_{\mathbf{F}}(S)$ coincides with $\overline{d}(S)$ as defined in Section 1 when $\mathbf{F}$ is the standard Følner sequence $\{1,\ldots,N\}$, $N=1,2,3,\ldots$
@@
Theorem 4.1. Let $\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ be Følner sequences in $\mathbb{Z}$. If $S_{1},\ldots,S_{k}\subseteq\mathbb{Z}$ satisfy $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$, then there exists $S\subseteq\mathbb{Z}$ such that $d(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1}(S)\subseteq D:=\bigcap_{i=1}^{k}\Delta_{2}(\mathbf{F}_{i},S_{i})$. Furthermore, there are $m_{1},\ldots,m_{\ell}\in\mathbb{Z}$, where
@@
One can also obtain an amplification of Theorem 3.2 similar to Theorem 4.1. For any function $g:\mathbb{N}\rightarrow[1,\infty)$ and any Følner sequence $\mathbf{F}$ in $\mathbb{N}$, let
@@
Theorem 4.2. Let $g:\mathbb{N}\rightarrow[1,\infty)$ and assume $([g(n)])_{n=1}^{\infty}$ is pointwise ergodic. Let $\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ be Følner sequences in $\mathbb{N}$. If $S_{1},\ldots,S_{k}\subseteq\mathbb{N}$ satisfy $\overline{d}_{\mathbf{F}_{i}}(S_{i})>0$ for each $i=1,\ldots,k$, then there exists $S\subseteq\mathbb{N}$ such that $d(S)\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$ and $\Delta_{1,g}(S)\subseteq D_{g}:=\bigcap_{i=1}^{k}\Delta_{2,g}(\mathbf{F}_{i},S_{i})$. Furthermore, we have that $\overline{d}(\Delta_{1,g}(S))\geq\prod_{i=1}^{k}\overline{d}_{\mathbf{F}_{i}}(S_{i})$.
We will not prove Theorem 4.1 or Theorem 4.1 here, as their proofs are practically the same as those of Theorems 1 and 3.2, respectively, with one modification which involves the use of a more general version of Theorem 2 in which $\overline{d}$ is replaced by $\overline{d}_{\mathbf{F}}$ (this result is a special case of Theorem 4.2 which is formulated in the next subsection). Theorem 4.1 is also a special case of Theorem 4.2, which will be proved in the next subsection.
Note that in Theorem 4.1, besides the Følner sequences $\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$, one more Følner sequence is implicitly present. Namely, $d(S)$ actually means $d_{\mathbf{F}}(S)$, where $\mathbf{F}$ is the standard Følner sequence $\{1,\ldots,N\}$, $N=1,2,3,\ldots$ The role of the Følner sequence $\mathbf{F}$ is different from the roles of the sequences $\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$. While $\mathbf{F}_{1},\ldots,\mathbf{F}_{k}$ are needed just to define appropriate notions of largeness for $S_{1},\ldots,S_{k}$, the role of the sequence $\mathbf{F}$ is to guarantee that the pertinent ergodic averages converge pointwise (which is instrumental in the proof of Theorem 2, formulated in Section 2). These ideas will be clarified in the next subsection where we extend Theorem 4.1 to the general amenable setup.
@@ 4.2 Extending Theorem 1 to amenable groups
As was mentioned in the introduction, a natural framework for generalizing Theorem 1 is that of amenable semigroups. In this subsection, we first introduce some definitions pertaining to the notion of amenability and establish an extension of Theorem 1 for amenable groups. While being of interest on its own, this result will serve as a tool for extending Theorem 1 to amenable semigroups, which will be done in subsection 4.3.
@@
Let $G$ be a countably infinite semigroup. A left Følner sequence in $G$ is a sequence $(F_{N})_{N=1}^{\infty}$ of finite non-empty subsets of $G$ satisfying
@@
for all $g\in G$. It is well-known that the existence of a left Følner sequence in a countably infinite semigroup $G$ implies that $G$ is left amenable (see [18, Theorem 6.4]). Furthermore, if $G$ is cancellative, [^9] A semigroup $G$ is (two-sided) cancellative if, for any $a,b,c\in G$, $ab=ac\Longrightarrow b=c$ and $ba=ca\Longrightarrow b=c$. then having a left Følner sequence is equivalent to $G$ being left amenable (see [2, Theorem 2]). From now on, we will tacitly assume that the semigroups we deal with are countable and cancellative.
Given a left Følner sequence $\mathbf{F}=(F_{N})_{N=1}^{\infty}$ in an amenable semigroup $G$ and a set $S\subseteq G$, define
@@
If $\overline{d}_{\mathbf{F}}(S)=\underline{d}_{\mathbf{F}}(S)$, then their common value is denoted by $d_{\mathbf{F}}(S)$.
For the remainder of the paper we will, as a rule, omit the adjective left when dealing with left amenable semigroups and left Følner sequences.
We will now embark on proving a version of Theorem 1 in the setup of amenable groups (Theorem 4.2 below). The further extension to amenable semigroups will be done in subsection 4.3.
Throughout the rest of this section, an important role is played by tempered Følner sequences.
Definition 4.3. For a group $G$, a sequence $(F_{n})_{n=1}^{\infty}$ of finite subsets of $G$ is called tempered if there is some constant $C>0$ such that
@@
for all $n\in\mathbb{N}$ with $n>1$. (For $A,B\subseteq G$ we define $A^{-1}B\vcentcolon=\{a^{-1}b:a\in A,b\in B\}$).
In much of what we do in this section, the following pointwise ergodic theorem, which utilizes the notion of a tempered Følner sequence, is essential.
Theorem 4.4 ([26], see Section 5.6 in [28]). Let $G$ be a group with tempered Følner sequence $(F_{N})_{N=1}^{\infty}$ and let $(X,\mathcal{B},\mu,(T_{g})_{g\in G})$ be a measure-perserving system. Then for any $f\in L^{2}$, the limit
$\lim_{N\rightarrow\infty}\frac{1}{|F_{N}|}\sum_{g\in F_{N}}f(T_{g}x)$
exists a.e.
In order to prove the generalization of Theorem 1 to countably infinite amenable groups we will need two results, Theorems 4.2 and 4.2, which can be viewed, correspondingly, as amenable analogs of Theorems 2 and 2.
Theorem 4.5. Let $G$ be a countably infinite amenable group with a tempered Følner sequence $\mathbf{F}=(F_{N})_{N=1}^{\infty}$. Let $(X,\mathcal{B},\mu,(T_{g})_{g\in G})$ be a measure-preserving system and let $B\in\mathcal{B}$ with $\mu(B)>0$. Then there is a set $P\subseteq G$ such that
$d_{\textbf{F}}(P)\geq\mu(B)\quad\text{and}\quad\mu\left(B\cap T_{g_{1}}^{-1}B\cap\cdots\cap T_{g_{r}}^{-1}B\right)>0\quad\text{for all}\quad g_{1},...,g_{r}\in P.$ (4.8)
Proof. After deleting (if needed) a set of measure zero from each $T^{-1}_{g}B$, we can and will assume without loss of generality that for any $g_{1},...,g_{j}\in G$, if $T^{-1}_{g_{1}}B\cap\cdots\cap T^{-1}_{g_{j}}B\neq\emptyset$, then $\mu(T^{-1}_{g_{1}}B\cap\cdots\cap T^{-1}_{g_{j}}B)>0$. For $N=1,2,3,\ldots$ define $f_{N}:X\rightarrow\mathbb{R}$ by,
@@
For all $N\in\mathbb{N}$ we have $\int f_{N}=\mu(B)$. Let $f=\limsup_{N\rightarrow\infty}f_{N}$. Note that $0\leq f_{N}\leq 1$ for all $N\in\mathbb{N}$, so $f$ is well-defined and takes values in $[0,1]$. Then by Fatou’s lemma, we have
$\int f=\int\limsup_{N\rightarrow\infty}f_{N}\geq\limsup_{N\rightarrow\infty}\int f_{N}=\mu(B).$
Since $f$ is nonnegative and $\int f\geq\mu(B)$, the set $f^{-1}[\mu(B),1]$ must have positive measure. By Theorem 4.2, $\lim_{N\to\infty}f_{N}$ exists a.e., so in particular there is some $x_{0}\in f^{-1}[\mu(B),1]$ such that $\lim_{N}f_{N}(x_{0})=f(x_{0})$. If we set $P^{\prime}\vcentcolon=\{g\in G:T_{g}x_{0}\in B\}$, then
$d_{\mathbf{F}}(P^{\prime})=\lim_{N\to\infty}\frac{|F_{N}\cap P^{\prime}|}{|F_{N}|}=\lim_{N\to\infty}\frac{1}{|F_{N}|}\sum_{g\in F_{N}}1_{B}({T_{g}}x_{0})=\lim_{N\to\infty}f_{N}(x_{0})=f(x_{0})\geq\mu(B).$
Now choose $h\in G$ so that $T_{h}x_{0}\in B$ and put $P=P^{\prime}h^{-1}$. We claim that $P$ satisfies formula (4.8). Firstly, observe $d_{\mathbf{F}}(P)=d_{\mathbf{F}}(P^{\prime})\geq\mu(B)$. Secondly, take $g_{1}h^{-1},\ldots,g_{r}h^{-1}\in P$, where $g_{1},\ldots,g_{r}\in P^{\prime}$. See that for $i=1,\ldots,r$ we have $T_{g_{i}}x_{0}\in B$, so $T_{g_{i}h^{-1}}T_{h}x_{0}\in B$, therefore $T_{h}x_{0}\in T_{g_{i}h^{-1}}^{-1}B$. Thus the intersection $B\cap T_{g_{1}h^{-1}}^{-1}B\cap\cdots\cap T_{g_{r}h^{-1}}^{-1}B$ is non-empty since it contains $T_{h}x_{0}$, so $\mu\left(B\cap T_{g_{1}h^{-1}}^{-1}B\cap\cdots\cap T_{g_{r}h^{-1}}^{-1}B\right)>0$. $\square$
For the proof of Theorem 4.2 we will need the following variant of Furstenberg’s correspondence principle in the amenable setup.
Theorem 4.6 ([7, Theorem 6.4.17]). Let $G$ be a countably infinite amenable group, let $\mathbf{F}$ be a Følner sequence in $G$, and let $S\subseteq G$ with $\overline{d}_{\mathbf{F}}(S)>0$. Then there is an invertible measure-preserving system $(X,\mathcal{B},\mu,(T_{g})_{g\in G})$ and a set $B\in\mathcal{B}$ with $\mu(B)=\overline{d}_{\mathbf{F}}(S)$ such that